OpenStudy (anonymous):

Please Please someone help. Last question to a 5 part assignment I've been trying to finish. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations. Graph of line going through ordered pairs (0, 300) and (450, 0)

OpenStudy (anonymous):

Find the slope by using rise over run then use point slope form

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

since x is higher on the 450 x2 becomes 450 x1 becomes 0 y2 become 300 y1 become 0 Slope is rise over run m=y2-y1/x2-x1 where m is the slope After you find slope, tell me

OpenStudy (anonymous):

300/450

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

0-300=-300 So the slope is -300/450 You have to reduce it so its -2/3 m=-2/3 The point slope form is y-y1=m(x-x1) or y-y2=m(x-x2) either one works then bring the y1 over the right side

OpenStudy (anonymous):

which one is y and which one is y1

OpenStudy (anonymous):

im confused

OpenStudy (nerdyotter):

(y1) is the value used for the y value you are replacing. So, if you had an example ordered pair of, (5,7), the equation would be [y-7 = m(x-5)], adding the slope to whatever equation you have.

OpenStudy (anonymous):

y-0=-2/3(x-450)

OpenStudy (nerdyotter):

So, if you understand rise over run, first find the slope of the two points, then use point slope form to substitute in your x, y, and slope values. An easy formula for finding slope in a two-dimensional coordinate plane is... \[\frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

distribute it

OpenStudy (anonymous):

what do i do next

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

distributing by multiplying the paranthese for example 6(a-b) 6a-6b

OpenStudy (anonymous):

y-0=-2/3x-450

OpenStudy (anonymous):

like that^

OpenStudy (anonymous):

no you multiply the -2/3 by 450 which equals -300 so its y=-(2/3)x+300

OpenStudy (anonymous):

ok and then what do i do?

OpenStudy (anonymous):

thankyou by the way for helping me

OpenStudy (anonymous):

That is the equation. You are done.

OpenStudy (anonymous):

wow that wasnt that hard. so in the future all i do is find the slope, then do that other formula and then distribute?